OpenStudy (anonymous):

If Leah is 6 years older than Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue?

OpenStudy (amistre64):

L= S+6 ; and J= 5+L; S+J+L=41. S=L-6; J=5+L; L=L. To wit; (L-6)+(5+L)+L=41 ->3L-1=41 -> 3L=42. L= 42/3 = 14 Leah is 14; Sue is 14-6 = 8; and John is 14+5=19. 14+8+19= 22+19 = 41 John should be moving out soon :)

OpenStudy (anonymous):

Sue is 8.

OpenStudy (anonymous):

take sue's age to be x. Sue= x Leah = x + 6 yrs old. John = x + 6 +5 yrs old. total = x + x + 6 + x + 6 + 5 =3x + 17 = 41 yrs old 3x = 41 - 17 = 24 x(Sue's age) = 24 / 3 = 8 yrs old. i hope i helped u with your work :)